HOW LONG DOES IT TAKE A BALANCE OF $100 TO REACH $2000 IF INTEREST IS 4% per year compounded quaterly?
Answer #1

PV =100
FV =2000
Rate per period =4%/4 =1%
FV =PV*(1+r)^n
2000 =100*1.04^n
applying log on both sides we get
n =log(20)/log(1.04) =76.38 quarter
Number of years =76.38/4 =19.10 years